### v1

Abstract
The Extreme Learning Machine (ELM) is a growing statistical technique widely applied to regression problems. In essence, ELMs are single-layer neural networks where the hidden layer weights are randomly sampled from a specific distribution, while the output layer weights are learned from the data. Two of the key challenges with this approach are the architecture design, specifically determining the optimal number of neurons in the hidden layer, and the method’s sensitivity to the random initialization of hidden layer weights.
This paper introduces a new and enhanced learning algorithm for regression tasks, the Effective Non-Random ELM (ENR-ELM), which simplifies the architecture design and eliminates the need for random hidden layer weight selection. The proposed method incorporates concepts from signal processing, such as basis functions and projections, into the ELM framework. We introduce two versions of the ENR-ELM: the approximated ENR-ELM and the incremental ENR-ELM. Experimental results on both synthetic and real datasets demonstrate that our method overcomes the problems of traditional ELM while maintaining comparable predictive performance.
keywords:
ELM, kernel methods, random feature learning, non-parametric regression
1 Introduction
ELM is an emerging statistical technique widely used for regression, classification, clustering, sparse approximation, feature learning, and compression. In this article, we focus on regression to simplify the discussion, but the idea presented can be adapted to other learning tasks. The most notable advantage of ELM is that there is no requirement for the hidden layer’s weights to be tuned. The hidden layer is randomly generated and is never updated after that. ELMs have been largely explored in the last decade; the reader can refer to [1] and [2] for interesting overviews. ELMs have been applied in many applicative contexts, but here we are more interested in the methodological point of view, therefore we do not assume any special conditions on the data. Furthermore, we would like to emphasize that the performance of an ELM compared to a traditional neural network with the same architecture has already been analyzed in the interesting work [3]. In [3] the authors show that there are cases in which ELMs have better accuracy than traditional neural networks, but above all ELMs are more convenient in terms of computation time, as they do not require the iterative back-propagation algorithm to learn weights. Our research aims to propose an alternative methodology for applying ELMs when they are preferred over other estimators, without claiming to propose a universal method that is suitable for every case. In this regard, we cite the same paper [3] “[Experimental results] confirm that for relatively small datasets, ELMs are better than network trained by the backpropagation algorithm”, meaning that ELMs are preferred when the dimension of the input space is relatively small, hence this will be our working assumption.
1.1 Motivation
The motivation for this paper stems from both our practical experience and a thorough review of the current literature. Providing a comprehensive summary of the existing research is challenging because many articles discuss ELMs and potential improvements. However, we believe that two recent papers [1] and [4] offer solid overviews of both the strengths and limitations of this approach, highlighting areas that require further attention. Based on our own studies, we have identified three key issues in the application of ELMs, which form the foundation of our research rationale.
Motivation (i) From the recent review [1] we cite “[ELM] still suffer from design choices, translated in free parameters, which are difficult to set optimally with the current mathematical framework, so practically they involve many trials and cross-validation to find a good projection space [number of hidden neurons].” The first motivation for our research is then the lack of an effective model selection criterion, or in other words, the lack of an effective architecture design.
Motivation (ii) In the same review [1], we cite “[ELM] … on top of the selection of the number of hidden [neurons] and the nonlinear [activation] function strongly depend on the particular random projection”. In [4], we cite “The supervisory mechanism reported in this paper clearly reveals that the assignment of random parameters of learner models should be data dependent to ensure the universal approximation property of the resulting randomized learner model.” One of the more interesting attempts at a solution to this issue is given by [5], where a data-dependent weights matrix for the hidden layer is proposed. Specifically in [5] a proper weights matrix that sorts the input data $\mathbb{x}_{i}$ by a well defined lexicographical order is established, the resulting hidden layer output matrix is proved to be non-singular, then, for the evaluation of the output weights, a classical OLS algorithm is applied. However, the proposal in [5] is specific for radial basis function activation and does not include model selection. The second motivation for our research is then to provide a data-dependent deterministic mechanism to fix the hidden layer output matrix, justifying the name “Non-Random” for our proposed method.
Motivation (iii) In [3], we cite “There are no smart algorithms for the inverse matrix calculation so that determining weights in the output layer for challenging dataset become feasible and memory efficient.” The third motivation for our research is then to furnish a very smart and efficient algorithm for determining weights in the output layer without employing matrix inversion.
1.2 Contribution
In this paper, we propose the Effective Non-Random ELM (ENR-ELM), a new method for non-parametric regression tasks. This method integrates signal processing concepts, such as basis functions and projections, into the ELM framework. The ENR-ELM algorithm operates in two phases: (1) the selection of a data-dependent hidden layer’s weights matrix and (2) the evaluation of the output layer weights.
In phase (1), the expected Gram matrix of the input data vectors $\mathbb{x}_{i}$, transformed by the hidden layer’s random feature map, is computed along with its eigen-decomposition. The orthonormal matrix $U$, whose columns are the eigenvectors of the Gram matrix, is then used to derive an optimal weights matrix $\hat{W}$, which represents the data-dependent selection of hidden layer’s weights.
Phase (2) can be implemented in two ways, resulting in two algorithms: the approximate ENR-ELM and the incremental ENR-ELM. In the first algorithm, the matrix $U$ is used as a proxy for the hidden layer output matrix, while in the second, the hidden layer output matrix is used in a classical incremental forward-stagewise regression. In both cases, the test error evaluation guides the selection of the optimal number of columns, corresponding to the appropriate number of neurons in the hidden layer.
To determine the optimal weights matrix $\hat{W}$ in phase (1) of the procedure, the network’s activation function must be invertible and take values in the range $[-1,1]$. Therefore, in the examples presented in this paper, we use the Erf activation function. The key distinction between our ENR-ELM and the traditional ELM lies in the choice of the hidden layer weight matrix, which is data-dependent rather than random in our approach. Although the spectral decomposition of the Gram matrix is computationally intensive, this cost is offset by the linear complexity of the second phase, making the overall procedure more efficient compared to the traditional ELM, particularly in terms of model selection.
1.3 Paper structure
The remainder of the paper is organized as follows: In Section 2, we review the Extreme Learning Machine (ELM) and the Neural Network Gaussian Process (NNGP) kernel, which arises when considering random neural networks in the infinite-width regime. Section 3 introduces the proposed ENR-ELM method and outlines its pseudo-algorithm. Extensive experimental results are presented in Section 4. Finally, Section 5 summarizes our conclusions and discusses future research directions.
2 Background
2.1 Notations
We denote the identity matrix of dimension $d$ by $I_{d}$. We use bold notation to denote a vector, which should be considered as a column vector unless stated otherwise. We denote matrices by capital letter and the $j$-th column of matrix $H$ by $H_{\cdot j}$, the $i$-th row of matrix $H$ by $H_{i\cdot}$; moreover if $J$ is a set of indices, then $H_{\cdot J}$ represents the submatrix of $H$ restricted to the columns with indices in $J$. The symbol $J^{c}$ denotes the complement of the set $J$. Finally, the following notation applies $\mathbb{1}_{n}=(1,\ldots,1)\in\mathbb{R}^{n}$.
2.2 ELM
Given a vector $\mathbb{x}\in\mathbb{R}^{n_{0}}$, its random projection is defined as
$z(\mathbb{x})=(z_{1}(\mathbb{x}),\ldots,z_{n}(\mathbb{x})):=\sigma(W\,\mathbb{x})\in\mathbb{R}^{n}$ (1)
where $W\in\mathbb{R}^{n\times n_{0}}$ is a random matrix with independent entries $W_{ij}\sim{\cal{N}}(0,1/\sqrt{n_{0}})$, and $\sigma:\mathbb{R}\to\mathbb{R}$ is referred to as the activation function, applied entry-wise to the vector $W\,\mathbb{x}$. Given the activation function $\sigma$, the weights matrix $W$, and a weights vector $\boldsymbol{\beta}\in\mathbb{R}^{n}$, a shallow fully connected neural network is the following function
$\mathbb{x}:=(x_{1},\ldots,x_{n_{0}})\in\mathbb{R}^{n_{0}}\mapsto f_{(\sigma,W,\boldsymbol{\beta})}(\mathbb{x}):=\boldsymbol{\beta}^{\top}\sigma(W\,\mathbb{x})\,\in\mathbb{R}.$
Consider a regression task with training sample $\{(\mathbb{x}_{i},y_{i})\}_{i=1,...,T}$, where $\mathbb{x}_{i}\in\mathbb{R}^{n_{0}}$ and $y_{i}\in\mathbb{R}$. The ELM first maps the input vector $\mathbb{x}$ into a high (theoretically even infinite) dimensional feature space, then applies a linear model into this feature space producing the following non-linear estimator
$\hat{f}_{(\sigma,W,\hat{\boldsymbol{\beta}})}(\mathbb{x})=\sum_{i=1}^{n}\hat{\beta}_{i}z_{i}(\mathbb{x})=\hat{\boldsymbol{\beta}}^{\top}\,\sigma(W\,\mathbb{x}),$ (2)
with $\hat{\boldsymbol{\beta}}$ learned from data by minimizing the following empirical risk
$\hat{\boldsymbol{\beta}}=\text{argmin}_{\boldsymbol{\beta}\in\mathbb{R}^{n}}\,\,\sum_{l=1}^{T}\left(y_{l}-\boldsymbol{\beta}^{\top}\sigma(W\,\mathbb{x}_{l})\right)^{2}.$ (3)
To clarify, ELM is a 1-layer (shallow) neural network where only the weights of the output layer are learned from the data, while the weights in the hidden layer are sampled from a Gaussian distribution and not tuned.
Consider data matrix $X=[\mathbb{x}_{1},\mathbb{x}_{2},\ldots,\mathbb{x}_{T}]\in\mathbb{R}^{n_{0}\times T}$, hidden layer output matrix $S=\sigma(WX)^{\top}=[\sigma(W\,\mathbb{x}_{1})^{\top},\,\,\sigma(W\,\mathbb{x}_{2})^{\top},\,\ldots\,,\sigma(W\,\mathbb{x}_{T})^{\top}]\in\mathbb{R}^{T\times n}$ and data vector $\mathbb{y}=[y_{1},y_{2},...,y_{T}]\in\mathbb{R}^{T}$, then in matrix notation problem in eq. (3) becomes
$\hat{\boldsymbol{\beta}}=\text{argmin}_{\boldsymbol{\beta}\in\mathbb{R}^{n}}\,\,\left\|\mathbb{y}-S\boldsymbol{\beta}\right\|_{2}^{2}$ (4)
a linear model with stochastic design matrix $S$.
If the matrix $S^{\top}\,S$ has full rank, then the solution is given by
$\hat{\boldsymbol{\beta}}=(S^{\top}\,S)^{-1}S^{\top}y,$ (5)
otherwise, the solution can be obtained by different approaches. It is possible to use the incremental forward-stagewise regression or penalized regression approaches such as Ridge regression or LASSO regression. In the case of Ridge regression the minimization given in eq. (4) is penalized by the $l_{2}$ norm of the regression coefficients and the solution is given by
$\hat{\boldsymbol{\beta}}=(S^{\top}\,S+\lambda I_{n})^{-1}S^{\top}y=S^{\top}(S\,S^{\top}+\lambda I_{T})^{-1}y,$ (6)
where the positive parameter $\lambda$ is chosen to avoid overfitting and preserve a good generalization property. Sometimes this choice is replaced by an early stopping criterion of the gradient descent method applied to minimization in eq. (4). While the incremental forward-stagewise regression or the LASSO regression guarantees variable selection (which is important for determining the relevants hidden layer neurons), the Ridge regression does not guarantee variable selection. In addition, we note that evaluating the expression in eq. (5) results in a computational cost that is cubic in the number of features, while evaluating the second expression in eq. (6) results in a computational cost that is cubic in the number of data points, which can be unfeasible in a data rich situation.
The choice of the number of features $n$ (i.e. the number of neurons in the hidden layer) is important and significantly depends on the structure of the input dataset $X$. As explained in the introduction, we do not consider any specific type of data, but our working hypothesis is that the input space dimension $n_{0}$ is relatively small. Therefore, we can think the input space to be an Euclidean space of limited dimension, and the random feature map (defined in eq. (1)) as a mechanism for scattering the data into a much higher-dimensional space, enhancing certain data characteristics. It is somewhat like the opposite of dimensionality reduction; here, the data dimensionality is increased to reveal hidden structures, allowing a linear operator to perform effectively in the new space. Mathematically, the data structure is captured by the Gram matrix. A significant area of research is understanding how the Gram matrix of the mapped features, $SS^{\top}\in\mathbb{R}^{T\times T}$, relates to the Gram matrix of the input features, $X^{\top}X\in\mathbb{R}^{T\times T}$. Some results in this area, specific to the case of ReLU activation function, are provided in [6]. The choice of the number $n$ of hidden neurons, which in this paper is referred to as model selection or architecture design, is a critical aspect; practitioners usually select $n$ by using the test error or some cross-validation criterion. One effective attempt to solve this problem is the incremental constructive approach proposed in [7], where a starting small-sized network is incrementally enlarged by generating hidden layer neurons and output weights until a predefined termination criterion is met.
As pointed out in motivation (i), there is still a lack of effective model selection criteria for ELM; our proposed method address this issue.
Moreover, as described in motivation (ii), beyond the choice of the number of hidden neurons $n$, ELM performance is conditioned by the particular sampling of the random matrix $W$; hence, some machine learning practitioners generate independent realizations of the ELM and subsequently apply some voting strategy, for example in [8]. Our proposed method accounts for this issue too, since the weights matrix $W$ will be learned from data and will not be stochastic.
Finally, as described in motivation (iii), matrix inversion in eq.  (5) or eq. (6) is computationally expensive if one considers that for each choice of $n$ and for each realization of $W$ it must be performed from scratch. Our proposed method also alleviates this issue because, apart from the eigen-decomposition of the Gram matrix, it does not require any matrix inversion to evaluate the coefficients in the last layer for different choices of $n$.
2.3 Neural Network Gaussian Process Kernel
In recent years, there has been a significant body of research focused on neural networks with stochastic weights in the infinite-width regime. Since the seminal work published in 1996 [9], multiple studies have shown that a neural network with independent Gaussian weights converges to a Gaussian process with zero mean and a known covariance matrix as the number of neurons in the hidden layers approaches infinity. While the theoretical understanding of this topic is quite extensive, see e.g. [10], [11], [12], the practical knowledge for using these results is still in its infancy. In this regard, a software package capable of calculating the limiting process covariance matrix, called Neural Network Gaussian Process (NNGP) kernel, has been made available to the scientific community for the first time in the paper [13]. For completeness, in the following we provide the definition of NNGP kernel given in [13] for a general deep neural network.
Consider a fully connected L-layered neural network $f:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}$ for $L\geq 1$ defined as
$f_{(\sigma,\mathcal{W})}(\mathbb{x})={\mathbb{w}^{(L+1)}}^{\top}z^{(L)}(\mathbb{x}),\quad z^{(l)}(\mathbb{x})=\sigma(W^{(l)}z^{(l-1)}(\mathbb{x})),\quad z^{(0)}(\mathbb{x})=\mathbb{x},$ (7)
where $\mathcal{W}:=vec(\mathbb{w}^{(L+1)},\cup_{l=1}^{L}W^{(l)})$, with $\mathbb{w}^{(L+1)}\in\mathbb{R}^{n_{L}}$ and $W^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ for $l\geq 1$, is the collection of all weight parameters. Note that the definition given in eq. (7) coincides with the definition given in eq. (2) when $L=1$ and $\mathcal{W}:=vec(\boldsymbol{\beta},W)$, being the weights of the output layer $\mathbb{w}^{(L+1)}=\boldsymbol{\beta}$ and weights of the hidden layer $W^{(1)}=W$.
In [13], for simplicity, a fully connected L-layered neural network with the same number of neurons in the hidden layer, i.e. $n_{l}=n$ for all $l=1,...,L+1$ is considered. This assumption is not necessary to achieve the convergence results presented in the literature which hold for $\text{min}\{n_{1},...,n_{L+1}\}\rightarrow\infty$, see [10], however, this hypothesis simplifies exposition.
Definition 1 (Definition NNGP Kernel).
Consider a stochastic fully connected L-layered neural network, as the one given in eq. (7) where all elements of $\mathcal{W}$ are independent Gaussian variables with zero mean and variance $1/n$, except in the first hidden layer where the variance is $1/n_{0}$, then the NNGP kernel is the covariance of the limiting Gaussian process defined as
$\Sigma_{\sigma}^{L+1}(\mathbb{x},\mathbb{y}):=plim_{n\rightarrow\infty}<f_{(\sigma,\mathcal{W})}(\mathbb{x}),f_{(\sigma,\mathcal{W})}(\mathbb{y})>,\quad\quad\forall\,\mathbb{x},\mathbb{y}\in\mathbb{R}^{n_{0}},$ (8)
where $plim$ stands for convergence in probability.
Note that in the original paper [13], the weights are all independent and identically distributed $\mathcal{N}(0,1)$ because in the definition of a fully-connected L-layered neural network given in eq. (7) the matrix of weights at layer $l$ is rescaled by the factor $1/\sqrt{n_{l-1}}$.
In [13] the authors proposed a fast algorithm that evaluates (for activation functions admitting exact dual activation, like erf,ReLU, etc) or approximates (for those activation functions not admitting exact dual activation) any L-layered NNGP kernel matrix. This algorithm is available at https://github.com/google/neural-tangents as a Python package; its manual furnish the complete list of activation functions for which to evaluate or approximate the NNGP kernel. The method described in [13] is based on the following recursive formula for the NNGP kernel: start from the input layer Kernel $K_{\sigma}^{(0)}(\mathbb{x},\mathbb{y}):=<\mathbb{x},\mathbb{y}>$ and define the successive kernels, for $l=1,...,L+1$, by
$K_{\sigma}^{(l)}(\mathbb{x},\mathbb{y}):=\mathbb{E}_{(u,v)\sim\mathcal{N}(0,\Lambda_{\sigma}^{(l)})}(\sigma(u)\sigma(v))$
where the covariance matrix is $\Lambda_{\sigma}^{(l)}:=\left[\begin{array}[]{cc}K_{\sigma}^{(l-1)}(\mathbb{x},\mathbb{x})&K_{\sigma}^{(l-1)}(\mathbb{x},\mathbb{y})\\ K_{\sigma}^{(l)}(\mathbb{x},\mathbb{y})&K_{\sigma}^{(l)}(\mathbb{y},\mathbb{y})\end{array}\right]\in\mathbb{R}^{2}$.
From our point of view, this package will be used to evaluate the NNPG kernel in the input data $\{\mathbb{x}_{i}\}_{i=1,...,T}$, which represents the Gram matrix of their transformation produced by the hidden layer of a stochastic shallow neural network with an infinite number of neurons.
3 The proposed method
Let us recall the well-known correspondence between deep neural networks and kernel methods, particularly the relationship between Extreme Learning Machines (ELMs) and kernel methods, as described in detail in [14].
In kernel methods, there is no need to explicitly define the feature map because the kernel, a non-negative definite bivariate function, is used to compute the Gram matrix of the transformed sample data and to generalize the estimator to new data. This ability to rely solely on the kernel, bypassing the explicit computation of the feature map, is referred to as the “kernel trick” in machine learning. On the other hand, in ELMs, the feature map — defined in eq. (1), which projects the input data $\mathbb{x}\in\mathbb{R}^{n_{0}}$ into a high-dimensional (of order $n$) or even infinite-dimensional space ($n\rightarrow\infty$) — is explicitly constructed, avoiding the use of the kernel. In both methods, the estimator is a linear combination of the features that have been transformed using the map and only the coefficients of this linear combination need to be learned from the data. The method proposed in this research combines these two approaches. As in kernel methods, the kernel is fixed, arguably the optimal choice you can have when using random feature map given in eq. (1); then we explicitly construct the optimal feature map, which determines the hidden layer output matrix of a shallow neural network. Subsequently, only the weights of the final layer need to be learned, as in ELMs.
We now have all the ingredients to present our proposed method. Consider a random shallow neural network with an infinite number of neurons in its hidden layer, and consider the NNGP kernel $K_{\sigma}=(K_{\sigma}^{(1)}(\mathbb{\mathbb{missing}}x_{i},\mathbb{\mathbb{missing}}x_{j}))_{i,j=1,...,T}\in\mathbb{R}^{T\times T}$ evaluated at the training input vectors by the exact method presented in Section 2. Matrix $K_{\sigma}$ represents the expected Gram matrix of the input data when transformed through the hidden layer of a random shallow neural network with an infinite number of neurons. Matrix $K_{\sigma}$ is known to be non-negative defined, let $K_{\sigma}=U\,\Delta\,U^{\top}$ be its spectral decomposition and rearrange the columns of matrix $U$ according to their informative content, that is, ordering them based on the quantity $|<y,U_{.j}>|$ arranged in decreasing order.
In this way, for any $n\leq T$, the first $n$ columns of the eigen-matrix $U$ represent an orthonormal basis of the best (most informative) subspace of dimension $n$ into which the input data is mapped by the hidden layer transformation when using $n$ neurons. Although we can immediately evaluate the training vector prediction, as $\hat{\mathbb{y}}=U\,U^{\top}\mathbb{y}$, to generalize the proposed estimator we need an oracle providing us the feature map $\sigma(W\mathbb{x})$ such that $\sigma(W\mathbb{x}_{i})=U_{i.}$. The oracle does not exist, but we can find an optimal matrix $\hat{W}\in\mathbb{R}^{n\times n_{0}}$ such that $\sigma(\hat{W}X)^{\top}$ is similar to matrix $U$ as much as possible. We propose to solve the following
$\hat{W}=argmin_{W\in\mathbb{R}^{T\times n_{0}}}\left\|\sigma(WX)^{\top}-U\right\|^{2}.$ (9)
Since the columns of $U$ are by construction orthonormal, their entries are within the range $[-1,1]$; assuming that the activation function is invertible and takes values in $[-1,1]$, we obtain the solution of equation (9) in a closed form, i.e. $\hat{W}=\sigma^{-1}(U^{\top})\,X^{\top}\,(XX^{\top})^{-1}$, where $\sigma^{-1}$ is intended to be applied entry-wise. After evaluating the optimal matrix $\hat{W}$ we have two approaches. In the first approach, which we call A-ENR-ELM (approximated ENR-ELM), we learn the output layer coefficients using the OLS to regress $y$ on the matrix $U$, which is a proxy for the matrix $\sigma(\hat{W}X)^{\top}$. This approach computes the final layer coefficients $\hat{\boldsymbol{\beta}}=U^{\top}y$ with linear computational cost. In the second approach, we learn the output layer coefficients using the classical incremental forward stagewise regression of $y$ on matrix $S=\sigma(\hat{W}X)^{\top}$, for that reason we call this second approach I-ENR-ELM (incremental ENR-ELM). In this second approach, each neuron (which corresponds to a column of matrix $S$) is partially added to the estimator incrementally until some convergence criterion is achieved.
Finally, knowing matrix $\hat{W}$ and vector $\hat{\boldsymbol{\beta}}$, obtained by one of the two approaches, it is possible to generalize the proposed estimator; indeed given a new input data $\mathbb{x}\in\mathbb{R}^{n_{0}}$, the prediction is given by $\hat{y}(\mathbb{x})=\hat{\boldsymbol{\beta}}^{\top}\,\sigma(\hat{W}\mathbb{x})$. The prediction applied to a set of test data, which was not used for learning $\hat{W}$ and $\hat{\boldsymbol{\beta}}$, gives the test error used for model selection.
3.1 Algorithm
In the following, we summarize the proposed method using pseudo-code. We divide the algorithm into two phases. In the first phase (algorithm 1), we compute the matrix $\hat{W}$, which depends on the set of input vectors and the activation function $\sigma(\cdot)$. In the second phase, we compute the output layer weights $\hat{\boldsymbol{\beta}}$ using two possible approaches. In the case of the A-ENR-ELM (algorithm 2), we use the matrix $U$ as a proxy for the matrix $\sigma(\hat{W}X)^{\top}$, and the evaluation of the output layer weights is straightforward. In the case of the I-ENR-ELM (algorithm 3), we use the matrix $S=\sigma(\hat{W}X)^{\top}$ and incrementally add small contributions from each neuron until the residual vector stops improving. Finally, for completeness, we also provide the algorithm for calculating the error curve (algorithm 4). When the error curve is evaluated on the test dataset, we obtain the test error curve that can be used to make model selection. Specifically, we propose to choose how many and which neurons to use in the hidden layer of the ENR-ELM by looking at the test error curve.
 algorithm 1 Evaluation of matrix $\hat{W}$
Require $\sigma:\mathbb{R}\to\mathbb{[}-1,1]$ invertible
Input: $X\in\mathbb{R}^{n_{0}\times T}$$\triangleright$ $X$ standardized
Evaluate NNGP kernel matrix $K_{\sigma}\in\mathbb{R}^{T\times T}$
Evaluate spectral decomposition $K_{\sigma}=U\,\Delta\,U^{\top}$
Evaluate matrix $\hat{W}=\sigma^{-1}(U^{\top})\,X^{\top}\,(XX^{\top})^{-1}\in\mathbb{R}^{T\times n_{0}}$
Return $\hat{W}$
 algorithm 2 Evaluation of vector $\hat{\boldsymbol{\beta}}$ for A-ENR-ELM:
Input $\mathbb{y}$, $U$, $n$, $\hat{W}$triangleright $y$ centered
$\triangleright$ $n$ maximum number of neurons
Initialize $\hat{J}=(\,)$
[…]
     select $j^{*}=\text{argmax}_{j\in\hat{J}^{c}}|<\mathbb{y},U_{.j}>|$
     update $\hat{J}=(\hat{J}\,,j^{*})$
end for
Evaluate $\hat{\boldsymbol{\beta}}=(U_{\cdot\hat{J}})^{\top}\mathbb{y}$
Update $\hat{W}=\hat{W}_{\cdot\hat{J}}$
Return $\hat{W}\,,\hat{\boldsymbol{\beta}}$
 algorithm 3 Evaluation of vector $\hat{\boldsymbol{\beta}}$ for I-ENR-ELM:
Input $y$, $S=\sigma(\hat{W}X)^{\top}$, $n$, $toll$, $\epsilon$, $\hat{W}$triangleright $y$ and $S$ centered
Initialize $\hat{\boldsymbol{\beta}}=0$, $\hat{\mathbb{y}}=\bar{\mathbb{y}}$, $\mathbb{r}=\mathbb{y}-\hat{\mathbb{y}}$, $\mathbb{r}_{old}=10^{10}\mathbb{1}$ and $\hat{J}=(\,)$
while #$\#(\hat{J})\leq n$ do
     select $j^{*}=\text{argmax}_{j}|<\mathbb{r},S_{.j}>|$
     update $\hat{\beta}_{j^{*}}=\hat{\beta}_{j^{*}}+\epsilon<r,S_{.j^{*}}>/\|S_{.j^{*}}\|^{2}$
     update $\hat{\mathbb{y}}=\hat{\mathbb{y}}+\epsilon<r,S_{.j^{*}}>/\|S_{.j^{*}}\|^{2}*S_{.j^{*}}$
     update $\mathbb{r}=\mathbb{r}-\epsilon<r,S_{.j^{*}}>/\|S_{.j^{*}}\|^{2}*S_{.j^{*}}$
     if $j^{*}\texttt{ not in }\hat{J}$ then $\hat{J}=(\hat{J}\,,j^{*})$
     if $(\|\mathbb{r}_{old}\|-\mathbb{\|}r\|)/\|\mathbb{y}\|<toll$ then break
     update $\mathbb{r}_{old}=\mathbb{r}$
end while
Update $\hat{W}=\hat{W}_{\cdot\hat{J}}$
Update $\hat{\boldsymbol{\beta}}=\hat{\boldsymbol{\beta}}_{\hat{J}}$
Return $\hat{W}\,,\hat{\boldsymbol{\beta}}$
 algorithm 4 Error curve evaluation
Input $\mathbb{y}$, $X$,$\hat{W}$, $\hat{\boldsymbol{\beta}}$, $\triangleright$ $y$ centered by the mean of the training
Evaluate $S=\sigma(\hat{W}X)^{\top}$$\triangleright$ $X$ standardized by the mean of and std of the training
Initialize $\hat{\mathbb{y}}=\bar{\mathbb{y}}$
[…]
     update $\hat{\mathbb{y}}=\hat{\mathbb{y}}+\hat{\beta}_{l}\,S_{.l}$
     evaluate error $Err(l)=\|\mathbb{y}-\hat{\mathbb{y}}\|/\|\mathbb{y}\|$
end for
Regarding the second phase of the proposed method, the choice of the parameter $n$, the maximum number of columns (i.e. the maximum number of neurons) depends on the dataset size. Notably, $n>>n_{0}$ since the dimension of the feature map’s target space is expected to exceed that of its input space. Similarly, in data-rich scenarios, where $T$ is significantly large, constructing error curves as the number of neurons approaches $T$ becomes unnecessary. It needs to be noted that both approaches assign importance to the columns of $\hat{W}^{\top}$ and rearrange them accordingly. Then, the first approach (A-ENR-ELM) considers neurons up to the maximum number $n$; while the second approach (I-ENR-ELM) saturates the number of neurons when the convergence criterion is met. In the respective algorithms, the ordered set $\hat{J}$ enables this rearrangement and selection mechanism.
Concerning the I-ENR-ELM, the choice of the parameter $\epsilon$ is critical, with $\epsilon\in(0,1]$. When $\epsilon=1$, the algorithm progresses with larger steps, following the classical forward stagewise regression. Conversely, when $\epsilon<1$, the steps are smaller, resulting in slower convergence. However, as noted in [15] cf Section 3.3.3 “this slow fitting can pay dividends in high-dimensional problems”.
We stress that the output parameter $\hat{W}$ and $\hat{\boldsymbol{\beta}}$ returned by the algorithm do not represent those of a single ENR-ELM, but rather represent the parameters of a nested family of ENR-ELMs that differ by one hidden layer neuron each. This is extremely important since it allows the construction of the test error curve, as explained in the algorithm 4, speeding up the entire model selection procedure. On the other hand, for traditional ELMs, the test error curve evaluation is performed by evaluating for each fixed number of neurons one or more realizations of the ELM, requiring a computational cost several orders of magnitude higher than the one we propose. This difference is highlighted by the experimental results presented in the next section.
4 Numerical study
In this section, we present numerical experiments performed on synthetic and real datasets to compare traditional ELM and the proposed ENR-ELM methods in terms of prediction accuracy and computational time. We ran all experiments on a computer with the specification given in Table 1. All the results are reproducible using the code available at https://github.com/FabianoVeglianti/Effective-Non-Random-Extreme-Learning-Machine. All experiments were implemented in Python 3.10; the required libraries are specified in the code repository.
Table 1: Specification of the computer on which the experiments were run
Parameter Value
Processor Intel i5-9300H
RAM size 8 GB
GPU NVidia GeForce GTX 1650
Operating system Windows 11 Home
[…]
4.1 Experimental setting
For each dataset considered in the experiments, we applied a random splitting into a training set (75% of the data) and a test set (25% of the data). To ensure comparability between features, we applied standardization to the training set by removing the mean and scaling the data to unit variance. The mean and the standard deviation of the training dataset were used to standardize the test set before evaluating the prediction accuracy.
We applied the algorithms proposed in Section 3.1 to the training data to obtain both the A-ENR-ELM model parameters and the I-ENR-ELM model parameters; after that, we used these parameters to compute training and test error curves by the procedure described in algorithm 4. The total computational time for the proposed ENR-ELM models includes the NNGP kernel matrix evaluation, its spectral decomposition, and the training and test error curves evaluation.
For the sake of comparison, we performed traditional ELM on the same datasets, and for each fixed number of hidden layer neurons, we evaluated 20 independent realizations of the ELM. For each run, we recorded the training and test errors and reported the mean of these values across the 20 runs. In terms of computational time, we measured the total time required for all 20 runs as well as the average time per individual run.
To account for fluctuations in the computational time measurements for model selection, which may be caused by background processes independent of human activities, the procedure was repeated 10 times. The mean elapsed time across these repetitions is reported to provide a more robust estimate.
The model selection process involves the evaluation of ELM, A-ENR-ELM, and I-ENR-ELM, with the number of hidden layer units set to a maximum value of $n=\min{\{50\times n_{0},\#Training/2\}}$. This choice is motivated by two considerations: first, it ensures that the hidden layer can project the data into a significantly larger space than the input space, facilitating a more expressive feature transformation; second, it guarantees that the resulting minimization given in eq. (4) remains well-conditioned, even for datasets with limited sample sizes, thus contributing to the numerical stability of the training process.
The training and test errors were measured with the normalized Root Mean Squared Error (RMSE) metric given in the following expression
$\|\mathbb{y}-\hat{\mathbb{y}}\|/\|\mathbb{y}\|=\sqrt{\sum_{i=1}^{T}(y_{i}-\hat{y}_{i})^{2}/\sum_{i=1}^{T}y_{i}^{2}}$ (10)
where $\mathbf{y}$ represents the true target values, $\mathbf{\hat{y}}$ denotes the predicted target values, and T is the number of samples. We normalize the RMSE by the L2-norm of the true target variable to enable comparisons across datasets with different target variable magnitudes. This ensures that the error magnitude remains independent of the target variable scale.
4.2 Results on synthetic datasets
Each synthetic dataset was created by generating $T\in\{300,1200\}$ pairs $\{(\mathbb{x}_{l},y_{l})\}_{l=1,...,T}$ where $\mathbb{x}_{l}\in\mathbb{R}^{n_{0}}$ with dimension $n_{0}=20$ or $n_{0}=80$ and $y_{l}=f(\mathbb{x}_{l})+\epsilon_{l}$, with $f(\cdot)$ being a given function and $\epsilon_{l}$ are independent and identically distributed Gaussian variables with zero mean and fixed variance $\sigma^{2}$. The generic input data vectors $\mathbb{x}$ were independently sampled from the following three different types of distribution: Uniform, where the vector components are independent and uniformly distributed in the interval $[-2\pi,2\pi]$; Indep. Gaussian, where the vector components are independent and identically distributed as a standard Gaussian variable; Toepl. Gaussian, where the vector $\mathbb{x}$ is distributed as $n_{0}$-variate Gaussian variable with zero mean and covariance given by $cov(x_{i},x_{j})=\rho^{|i-j|}$ with $\rho=0.8$. We consider two types of function $f(\cdot)$: Linear, where. $f(\mathbb{x})=\mathbb{\alpha}^{\top}\mathbb{x}+\beta$ where the components of $\mathbb{\alpha}$ and $\beta$ are independent and uniformly distributed in the interval $[-2,2]$, and Shallow NN, where $f(\mathbb{x})$ is the output obtained by a Neural Network with one hidden layer composed of 100 neurons whose hidden layer weights are drawn from a Gaussian distribution with zero mean and variance $1/n_{0}$, while the output layer weights are generated according to a Gaussian distribution with zero mean and variance $1/100$. Finally, we fix the variance of the noise $\sigma^{2}$ such that the Signal to Noise Ratio (SNR), defined as the ratio between the variance of the unknown signal and the variance of the noise, results to be $SNR=2$ or $SNR=10$, the first case representing a strong noise case, the second case representing a weak noise case.
With the two different choices of the sample size ($T=300$ or $T=1200$), two different choices of the input dimension ($n_{0}=20$ or $n_{0}=80$), the three different choices of input data generation (Uniform or Indep. Gaussian or Toepl. Gaussian), the two different choices of unknown function (Linear, Shallow NN) and the two different choices of SNR ($SNR=2$ or $SNR=10$) we have a total of 48 different datasets, which are summarized in Table 2.
Table 2: Characteristics of synthetic datasets
Dataset $T$ $n_{0}$ Input distribution $f(\cdot)$ SNR
Dataset 1 300 20 Uniform Linear 2
Dataset 2 300 20 Uniform Linear 10
Dataset 3 300 20 Indep. Gaussian Linear 2
Dataset 4 300 20 Indep. Gaussian Linear 10
Dataset 5 300 20 Toepl. Gaussian Linear 2
Dataset 6 300 20 Toepl. Gaussian Linear 10
Dataset 7 300 20 Uniform Shallow NN 2
Dataset 8 300 20 Uniform Shallow NN 10
Dataset 9 300 20 Indep. Gaussian Shallow NN 2
Dataset 10 300 20 Indep. Gaussian Shallow NN 10
Dataset 11 300 20 Toepl. Gaussian Shallow NN 2
Dataset 12 300 20 Toepl. Gaussian Shallow NN 10
Dataset 13 300 80 Uniform Linear 2
Dataset 14 300 80 Uniform Linear 10
Dataset 15 300 80 Indep. Gaussian Linear 2
Dataset 16 300 80 Indep. Gaussian Linear 10
Dataset 17 300 80 Toepl. Gaussian Linear 2
Dataset 18 300 80 Toepl. Gaussian Linear 10
Dataset 19 300 80 Uniform Shallow NN 2
Dataset 20 300 80 Uniform Shallow NN 10
Dataset 21 300 80 Indep. Gaussian Shallow NN 2
Dataset 22 300 80 Indep. Gaussian Shallow NN 10
Dataset 23 300 80 Toepl. Gaussian Shallow NN 2
Dataset 24 300 80 Toepl. Gaussian Shallow NN 10
Dataset 25 1200 20 Uniform Linear 2
Dataset 26 1200 20 Uniform Linear 10
Dataset 27 1200 20 Indep. Gaussian Linear 2
Dataset 28 1200 20 Indep. Gaussian Linear 10
Dataset 29 1200 20 Toepl. Gaussian Linear 2
Dataset 30 1200 20 Toepl. Gaussian Linear 10
Dataset 31 1200 20 Uniform Shallow NN 2
Dataset 32 1200 20 Uniform Shallow NN 10
Dataset 33 1200 20 Indep. Gaussian Shallow NN 2
Dataset 34 1200 20 Indep. Gaussian Shallow NN 10
Dataset 35 1200 20 Toepl. Gaussian Shallow NN 2
Dataset 36 1200 20 Toepl. Gaussian Shallow NN 10
Dataset 37 1200 80 Uniform Linear 2
Dataset 38 1200 80 Uniform Linear 10
Dataset 39 1200 80 Indep. Gaussian Linear 2
Dataset 40 1200 80 Indep. Gaussian Linear 10
Dataset 41 1200 80 Toepl. Gaussian Linear 2
Dataset 42 1200 80 Toepl. Gaussian Linear 10
Dataset 43 1200 80 Uniform Shallow NN 2
Dataset 44 1200 80 Uniform Shallow NN 10
Dataset 45 1200 80 Indep. Gaussian Shallow NN 2
Dataset 46 1200 80 Indep. Gaussian Shallow NN 10
Dataset 47 1200 80 Toepl. Gaussian Shallow NN 2
Dataset 48 1200 80 Toepl. Gaussian Shallow NN 10
[…]
Fig. 1 shows the training error of the ELM, A-ENR-ELM, and I-ENR-ELM models across the first 12 synthetic datasets with sample size $T=300$ and input dimension $n_{0}=20$ in terms of normalized root mean square error (RMSE). These results indicate that the proposed models avoid overfitting effectively. Specifically, for the A-ENR-ELM model, the training error saturates beyond a certain point, reflecting its ability to prevent overfitting. Similarly, the I-ENR-ELM model shows a plateau in training error upon reaching its stopping criterion, thereby preventing the addition of unnecessary neurons. By contrast, the traditional ELM model tends to fit all available data, potentially leading to interpolation and overfitting.
Fig. 2 presents the test error curves for the same models. This figure further corroborates the tendency of the traditional ELM to overfit the data, whereas both the A-ENR-ELM and I-ENR-ELM models demonstrate resilience against overfitting. Moreover, Fig. 2 highlights that the proposed models converge more rapidly to the minimum test error, and achieve performance comparable to that of the traditional ELM. In Fig. 1 and Fig. 2 the I-ENR-ELM curve transitions from solid to dashed at the point where the convergence criterion is met. Beyond this point, the last error value is replicated and shown as a dashed line to facilitate comparison with the other curves.
Table 3 presents the minimum observed test error achieved by each model across the first 12 synthetic datasets and the associated number of hidden layer units required. The results reveal that the I-ENR-ELM and traditional ELM models exhibit similar hidden layer neuron requirements. However, the A-ENR-ELM model generally demands more hidden units to achieve its minimum test error. Notably, even though the A-ENR-ELM model is highly resistant to overfitting, it continues to progressively fit the information in the data as the number of neurons increases. This continued adaptation leads to a steady decrease in training error, although marginally, as more neurons are added. This behavior causes the test error to fluctuate around a minimum value, typically reached early in the training process. As a result, using the minimum test error to determine the optimal number of neurons for the A-ENR-ELM model is not a reliable criterion, as this approach often yields an inflated estimate of the required number of neurons due to these test error fluctuations. A more appropriate approach for this model could be choosing the point at which the test error curve has the greatest curvature. Selecting the point where the test error curve shows the greatest curvature is generally the most suitable approach for all methods. Indeed, examining the shapes of the test error curves, we see that the minimum point reported in Table 3 is less parsimonious than the point of maximum curvature identifiable in Fig. 2.
So far, we have used only 12 of the 48 synthetic datasets. Similar conclusions can be drawn by analyzing the results obtained in the remaining synthetic datasets, see appendix A.
Table 4 contains the computational time required for all considered methods and 48 synthetic datasets. The results shown in Table 4 demonstrate that the proposed methods exhibit greater computational efficiency compared to the traditional ELM. Notably, the A-ENR-ELM proves to be significantly more computationally efficient. This comparison also highlights the key factors influencing the computational time of ENR-ELM algorithms: sample size and input dimensionality. Specifically, sample size predominantly determines the computational time for A-ENR-ELM, as observed from the almost constant computational time for datasets with sample size $T=300$, followed by a marked increase for larger datasets with sample size $T=1200$. This indicates that, for A-ENR-ELM, the primary computational cost arises from the eigen decomposition of the matrix $K_{\sigma}$. While this cost is also incurred by the I-ENR-ELM model, for this model also input dimensionality plays a prominent role. Finally, input dimensionality seems to be almost negligible for the time required by both A-ENR-ELM and ELM models. The results further indicate that the computational time for an exhaustive model selection using the traditional ELM algorithm increases with larger sample sizes. This increase can be attributed to the exhaustive nature of the model selection process, where the performance is evaluated across a range of hidden layer units from 1 to $n=\min{\{50\times n_{0},\#Training/2\}}$. Although a reduced number of hidden units could be considered for ELM to expedite the model selection process, doing so would introduce bias in the comparison, favoring ELM in terms of computational time while simultaneously disadvantaging it in terms of predictive accuracy. Therefore, maintaining consistency in model selection ensures a fair and comprehensive comparison.
Figure 1: Training error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with input dimension $n_{0}=20$. The ELM curve represents the mean training error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Figure 2: Test error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with input dimension $n_{0}=20$. The ELM curve represents the mean test error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Table 3: Optimal number of hidden layer neurons and relative test error (the minimum of the test error curve) for synthetic datasets with sample size $T=300$ and input dimension $n_{0}=20$. For the traditional ELM the minimum is evaluated on the mean test error curve across 20 realizations and its standard deviation is reported.
dataset A-ENR-ELM I-ENR-ELM ELM
n value n value n value
Dataset 1 15 0.5738 13 0.5741 24 0.5805 $\pm$ 0.0171
Dataset 2 32 0.3509 20 0.3562 30 0.3612 $\pm$ 0.0104
Dataset 3 85 0.5898 18 0.5976 21 0.5996 $\pm$ 0.0170
Dataset 4 61 0.2849 20 0.2871 26 0.2939 $\pm$ 0.0134
Dataset 5 51 0.5463 13 0.5511 23 0.5799 $\pm$ 0.0179
Dataset 6 76 0.3568 12 0.3439 22 0.3616 $\pm$ 0.0083
Dataset 7 75 0.6379 13 0.6418 21 0.6509 $\pm$ 0.0107
Dataset 8 39 0.4526 20 0.4579 22 0.4606 $\pm$ 0.0172
Dataset 9 10 0.6732 10 0.6732 22 0.6973 $\pm$ 0.0134
Dataset 10 79 0.4021 19 0.3991 31 0.4104 $\pm$ 0.0184
Dataset 11 92 0.5298 18 0.5362 19 0.5565 $\pm$ 0.0192
Dataset 12 96 0.3775 18 0.3761 27 0.3770 $\pm$ 0.0097
Table 4: Computational time on synthetic datasets expressed in milliseconds.
Dataset A-ENR-ELM I-ENR-ELM ELM (20 iterations) ELM (1 iteration)
Dataset 1 16.32146 124.96869 5307.37014 265.36851
Dataset 2 14.18106 140.60400 3787.67999 189.38400
Dataset 3 14.54270 117.02895 3738.63963 186.93198
Dataset 4 14.25233 121.38743 4370.76972 218.53849
Dataset 5 14.08346 105.09719 3819.09180 190.95459
Dataset 6 14.75797 107.49984 3645.48390 182.27419
Dataset 7 15.93076 115.34181 3760.09064 188.00453
Dataset 8 13.86249 114.22304 3777.84259 188.89213
Dataset 9 14.53062 94.65098 3748.16894 187.40845
Dataset 10 13.96850 143.14260 3757.50901 187.87545
Dataset 11 14.24737 106.52110 6491.89338 324.59467
Dataset 12 14.36173 136.39106 5169.10815 258.45541
Dataset 13 20.54475 333.03524 6069.08772 303.45439
Dataset 14 24.03600 435.90489 5367.00658 268.35033
Dataset 15 16.05688 335.69949 4970.70184 248.53509
Dataset 16 16.03769 408.20349 4678.19782 233.90989
Dataset 17 15.10580 278.86309 4474.52763 223.72638
Dataset 18 18.08123 403.01774 4430.76436 221.53822
Dataset 19 15.46903 312.33403 4566.51599 228.32580
Dataset 20 15.08143 369.08719 4375.69521 218.78476
Dataset 21 17.08805 329.02546 4371.82896 218.59145
Dataset 22 16.03508 370.74961 4475.03478 223.75174
Dataset 23 15.17585 262.05518 4395.86727 219.79336
Dataset 24 15.59896 300.90870 4304.27108 215.21355
Dataset 25 338.16093 687.32146 135352.25036 6767.61252
Dataset 26 325.93514 748.87876 130395.48539 6519.77427
Dataset 27 335.48370 764.41084 130706.77772 6535.33889
Dataset 28 329.64922 715.07219 131380.34171 6569.01709
Dataset 29 321.74117 595.92242 132626.76983 6631.33849
Dataset 30 326.42065 730.79328 128876.64502 6443.83225
Dataset 31 331.72239 666.00263 129583.91208 6479.19560
Dataset 32 321.78727 692.16751 130419.67952 6520.98398
Dataset 33 329.06390 717.21019 130369.22295 6518.46115
Dataset 34 328.76441 697.19879 131257.85343 6562.89267
Dataset 35 330.05414 579.63691 132284.08713 6614.20436
Dataset 36 325.93300 694.37587 150845.70616 7542.28531
Dataset 37 357.21692 1448.53690 140409.61809 7020.48090
Dataset 38 345.47655 1598.43018 138844.50821 6942.22541
Dataset 39 344.23160 1365.46915 140113.51526 7005.67576
Dataset 40 349.53607 1652.66787 142978.33630 7148.91681
Dataset 41 342.46273 1050.00356 138376.15035 6918.80752
Dataset 42 344.69478 1194.74726 139728.52306 6986.42615
Dataset 43 347.03367 1358.38221 137371.92784 6868.59639
Dataset 44 350.87913 1350.14249 141273.82691 7063.69135
Dataset 45 346.11636 1254.28499 141724.36133 7086.21807
Dataset 46 358.33137 1608.32152 138566.31555 6928.31578
Dataset 47 342.84621 1152.46139 140474.07547 7023.70377
Dataset 48 353.20407 1135.06300 139754.34675 6987.71734
[…]
4.3 Results on real datasets
The datasets used in this section were acquired from the UCI [16] repository, from the dataset repository of The Elements of Statistical Learning[15] [^1] https://hastie.su.domains/ElemStatLearn/data.html and from the University of Porto repository [^2] https://www.dcc.fc.up.pt/ ltorgo/Regression/. Table 5 contains the characteristics of the datasets along with their description.
Table 5: Characteristics and description of real datasets. The column ”number of features” ($n_{0}$) represents the total feature count after applying one-hot encoding to categorical variables. The code used for this encoding is available in the code repository.
Dataset $T$ $n_{0}$ Min. value of $y$ Max. value of $y$ Description
Abalone 4177 10 1.0 29.0 Predicting the age of abalone from physical measurements.
Auto MPG 392 21 9.0 46.6 The data concerns city-cycle fuel consumption in miles per gallon, to be predicted in terms of 3 multivalued discrete and 5 continuous attributes.
California Housing 20640 8 0.15 5.0 This dataset was derived from the 1990 U.S. census, using one row per census block group. The goal is to predict the median house value.
Delta Ailerons 7129 5 -0.0021 0.0022 Prediction of variance of ailerons positioning in the F-16 fighter aircraft.
LA Ozone 330 9 1.0 38.0 These data record the level of atmospheric ozone concentration from eight daily meteorological measurements made in the Los Angeles basin in 1976. The goal is to predict the upland maximum ozone.
Machine CPU 209 6 6.0 1550.0 Relative CPU Performance Data, described in terms of its cycle time, memory size, etc. The goal is to predict the CPU performance.
Prostate Cancer 97 8 -0.4308 5.5829 Clinical data for men who have undergone a prostatectomy. The goal is to predict the logarithm of the prostate-specific antigen (PSA) level.
Servo 167 19 0.1312 7.1 Predicting the rise time of a servomechanism in terms of two (continuous) gain settings and two (discrete) choices of mechanical linkages.
[…]
Fig. 3 illustrates the training error of the ELM, A-ENR-ELM, and I-ENR-ELM models across real datasets. The results on real datasets align with those observed on synthetic datasets: both A-ENR-ELM and I-ENR-ELM successfully avoid overfitting, whereas the traditional ELM exhibits overfitting. Additionally, the results reveal that for the Auto MPG dataset, the A-ENR-ELM model exhibits a higher saturation point in training error compared to I-ENR-ELM, suggesting that these two models may perform differently depending on the dataset.
Fig. 4 depicts the test error curves of the ELM, A-ENR-ELM, and I-ENR-ELM models across real datasets. This figure further confirms that the proposed methods converge more rapidly to the minimum error than the traditional ELM. In terms of overall predictive performance, the ENR-ELM methods demonstrate comparable generalization capabilities to the traditional ELM, with performance varying slightly across datasets. However, consistently with the observation on the training error curve, in the case of the Auto MPG dataset, the A-ENR-ELM model performs differently from the I-ENR-ELM model.
Table 6 shows the minimum test error achieved by each model across all real datasets, along with the number of hidden layer units required to reach this minimum. Unlike the results obtained on synthetic datasets, the I-ENR-ELM model attains the minimum error using fewer hidden units than the traditional ELM, highlighting its greater parsimony. On the other hand, as observed on synthetic datasets, the A-ENR-ELM model requires a greater number of hidden layer units to achieve the minimum error. Again, this is because the best criterion for the A-ENR-ELM model should be selecting the point of greatest curvature in the test error curve rather than its minimum point.
Table 7 contains the computational time required for model selection. The results support the findings from experiments on synthetic datasets. Notably, it reveals that the proposed models offer no substantial computational efficiency benefit over a single instance of the traditional ELM when the training set is large. In particular, when $T>5000$, as with the California Housing and Delta Ailerons datasets, there is even a decrease in efficiency. This is largely due to the high computational cost associated with the eigen-decomposition of the matrix $K_{\sigma}\in\mathbb{R}^{T\times T}$. In principle, this cost could potentially be mitigated by employing subsampling of the training dataset.
Figure 3: Training error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on real datasets. The ELM curve represents the mean training error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Figure 4: Test error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on real datasets. The ELM curve represents the mean test error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Table 6: Optimal number of hidden layer neurons and relative test error (the minimum of the test error curve). For the traditional ELM the minimum is evaluated on the mean test error curve across 20 realizations and its std is reported.
dataset A-ENR-ELM I-ENR-ELM ELM
n value n value n value
Abalone 51 0.6839 6 0.6878 29 0.6721 $\pm$ 0.0035
Auto MPG 14 0.4422 21 0.3789 26 0.3771 $\pm$ 0.0120
California Housing 387 0.6080 7 0.6160 160 0.5653 $\pm$ 0.0050
Delta Ailerons 248 0.5608 5 0.5622 53 0.5501 $\pm$ 0.0052
LA Ozone 4 0.5231 5 0.5269 14 0.5350 $\pm$ 0.0129
Machine CPU 25 0.3751 2 0.3543 25 0.3899 $\pm$ 0.0764
Prostate Cancer 24 0.6465 6 0.6530 6 0.6729 $\pm$ 0.0456
Servo 47 0.5031 10 0.4685 19 0.4860 $\pm$ 0.0203
Table 7: Computational time on real datasets expressed in milliseconds.
Dataset A-ENR-ELM I-ENR-ELM ELM (20 iterations) ELM (1 iteration)
Abalone $1.03\times 10^{4}$ $1.57\times 10^{4}$ $6.45\times 10^{5}$ $3.22\times 10^{4}$
Auto MPG $2.70\times 10^{1}$ $1.98\times 10^{2}$ $9.93\times 10^{3}$ $4.96\times 10^{2}$
California Housing $9.68\times 10^{5}$ $2.85\times 10^{6}$ $1.75\times 10^{6}$ $8.76\times 10^{4}$
Delta Ailerons $4.33\times 10^{4}$ $2.01\times 10^{5}$ $2.59\times 10^{5}$ $1.30\times 10^{4}$
LA Ozone $1.95\times 10^{1}$ $6.21\times 10^{1}$ $7.07\times 10^{3}$ $3.54\times 10^{2}$
Machine CPU $6.74\times 10^{0}$ $4.79\times 10^{1}$ $2.22\times 10^{3}$ $1.11\times 10^{2}$
Prostate Cancer $3.29\times 10^{0}$ $1.30\times 10^{1}$ $2.34\times 10^{2}$ $1.17\times 10^{1}$
Servo $6.24\times 10^{0}$ $8.03\times 10^{1}$ $1.16\times 10^{3}$ $5.82\times 10^{1}$
[…]
5 Conclusion
In this work, we introduced a novel method for non-parametric regression tasks called the Effective Non-Random Extreme Learning Machine (ENR-ELM), designed to address key issues present in traditional Extreme Learning Machines (ELM). Our approach selects the hidden layer weights in a data-dependent manner, which facilitates model selection, overcomes challenges associated with the random weight initialization of conventional ELMs, and eliminates the need for computationally expensive inverse matrix calculations to determine the output layer weights.
Our research followed a recent trend in literature involving a data-dependent hidden layer transformation. Specifically, our approach aims to generate the hidden layer features such that they approximate an orthonormal basis for the feature space generated by an infinite number of hidden layer neurons. We derive this target basis using the eigen-decomposition of the Gram matrix of such infinitely wide hidden layer transformation.
With this data-dependent hidden layer transformation in place, we proposed two distinct approaches for utilizing the generated features. The “approximated” approach operates under the assumption that the transformed hidden layer features perfectly match the desired orthonormal basis. Although this assumption is not strictly accurate, the method remains effective, thus the “approximated” designation. In contrast, the “incremental” approach relies solely on the actual features produced by the hidden layer transformation. Since the resulting design matrix for the Ordinary Least Squares problem in the output layer may not be full rank, we apply an iterative algorithm that adaptively weights each feature’s contribution.
Our empirical evaluation demonstrated that the proposed ENR-ELM significantly reduces the overall computational time required for both model selection and training, without sacrificing predictive performance. By introducing this data-dependent method for selecting hidden layer weights, we advance the field of neural network architectures that do not depend on backpropagation. While backpropagation remains the standard for achieving highly effective results, it is also computationally intensive. Our method provides an efficient alternative, highlighting the potential of non-backpropagation-based approaches.
Extreme learning itself is not necessarily limited to the 1-layer architecture; it can be extended to any type of MLP (MultiLayer Perceptron). In any case, the only weights learned from the data are those of the output layer, while all the weights of the hidden layers are left untrained. In this way, the model in eq. (2) for the regression function remains the same, the random projection $z(\mathbb{x})$ defined in eq. (1) are instead obtained as a concatenation of all the hidden layers. Since the addition of intermediate layers does not entail any difference in the mathematical treatment, we decided to present only ELM for simplicity of presentation. Finally, though ELM unifies classification and regression tasks, we only focused on regression. However we do not see particular difficulty in extending our proposed method to the classification task, apart from substituting the $l_{2}$ loss in eq. (3) with a class separability loss.
Declarations
  • Funding: Daniela De Canditiis was partially funded by ”INdAM - GNCS Project”, codice CUP-E53C23001670001
  • The authors declare no conflict of interest
  • Code availability: the code is available at https://github.com/FabianoVeglianti/Effective-Non-Random- Extreme-Learning-Machine.
Appendix A Additional Results
This section provides additional plots and tables for the remaining synthetic data-sets.
Figure 5: Training error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with sample size $T=300$ and input dimension $n_{0}=80$. The ELM curve represents the mean training error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Figure 6: Training error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with sample size $T=1200$ and input dimension $n_{0}=20$. The ELM curve represents the mean training error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Figure 7: Training error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with with sample size $T=1200$ and input dimension $n_{0}=80$. The ELM curve represents the mean training error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Figure 8: Test error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with with sample size $T=300$ input dimension of $n_{0}=80$. The ELM curve represents the mean test error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Figure 9: Test error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with sample size $T=1200$ input dimension of $n_{0}=20$. The ELM curve represents the mean test error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Figure 10: Test error of ELM, incremental-ENR-ELM, and approximated-ENR-ELM on synthetic datasets with with sample size $T=1200$ input dimension of $n_{0}=80$. The ELM curve represents the mean test error over 20 independent realizations of the ELM model, with the shaded region indicating the range between the minimum and maximum values observed across these realizations.
Table 8: Optimal number of hidden layer neurons and relative test error (the minimum of the test error curve) for synthetic datasets with sample size $T=300$ and input dimension $n_{0}=80$. For the traditional ELM, the minimum is evaluated on the mean test error curve across 20 realizations and its standard deviation is reported.
dataset A-ENR-ELM I-ENR-ELM ELM
n value n value n value
Dataset 13 78 0.6902 59 0.6844 83 0.6969 $\pm$ 0.0136
Dataset 14 66 0.3728 58 0.3659 90 0.4068 $\pm$ 0.0118
Dataset 15 68 0.7294 54 0.7125 79 0.7704 $\pm$ 0.0164
Dataset 16 52 0.4669 57 0.4527 80 0.4778 $\pm$ 0.0123
Dataset 17 13 0.6360 13 0.6594 52 0.6649 $\pm$ 0.0273
Dataset 18 28 0.3448 69 0.3433 80 0.3491 $\pm$ 0.0123
Dataset 19 87 0.8214 65 0.7903 82 0.8306 $\pm$ 0.0214
Dataset 20 90 0.5840 52 0.5906 83 0.6466 $\pm$ 0.0114
Dataset 21 106 0.7112 54 0.7005 76 0.7160 $\pm$ 0.0303
Dataset 22 111 0.3656 75 0.3661 83 0.3619 $\pm$ 0.0130
Dataset 23 24 0.6508 72 0.6238 81 0.6419 $\pm$ 0.0131
Dataset 24 99 0.3770 59 0.3649 83 0.3829 $\pm$ 0.0111
Table 9: Optimal number of hidden layer neurons and relative test error (the minimum of the test error curve) for synthetic datasets with sample size $T=1200$ and input dimension $n_{0}=20$. For the traditional ELM the minimum is evaluated on the mean test error curve across 20 realizations and its standard deviation is reported.
dataset A-ENR-ELM I-ENR-ELM ELM
n value n value n value
Dataset 25 20 0.6003 17 0.5977 27 0.6083 $\pm$ 0.0032
Dataset 26 247 0.3276 20 0.3304 30 0.3307 $\pm$ 0.0031
Dataset 27 269 0.5687 16 0.5729 25 0.5727 $\pm$ 0.0032
Dataset 28 397 0.3171 19 0.3233 31 0.3207 $\pm$ 0.0025
Dataset 29 78 0.5731 17 0.5776 30 0.5839 $\pm$ 0.0048
Dataset 30 21 0.3038 15 0.3001 34 0.3108 $\pm$ 0.0026
Dataset 31 439 0.6399 15 0.6415 26 0.6422 $\pm$ 0.0042
Dataset 32 40 0.4796 17 0.4842 26 0.4829 $\pm$ 0.0035
Dataset 33 233 0.6389 18 0.6412 25 0.6412 $\pm$ 0.0030
Dataset 34 157 0.3478 19 0.3502 33 0.3510 $\pm$ 0.0029
Dataset 35 138 0.6113 18 0.6150 26 0.6128 $\pm$ 0.0043
Dataset 36 362 0.3579 15 0.3630 49 0.3498 $\pm$ 0.0047
Table 10: Optimal number of hidden layer neurons and relative test error (the minimum of the test error curve) for synthetic datasets with sample size $T=1200$ and input dimension $n_{0}=80$. For the traditional ELM the minimum is evaluated on the mean test error curve across 20 realizations and its standard deviation is reported.
dataset A-ENR-ELM I-ENR-ELM ELM
n value n value n value
Dataset 37 202 0.6208 66 0.6242 84 0.6245 $\pm$ 0.0037
Dataset 38 190 0.3132 72 0.3131 101 0.3163 $\pm$ 0.0029
Dataset 39 449 0.5847 71 0.5830 85 0.5862 $\pm$ 0.0029
Dataset 40 265 0.3020 70 0.3108 87 0.3048 $\pm$ 0.0021
Dataset 41 389 0.5705 60 0.5782 82 0.5751 $\pm$ 0.0037
Dataset 42 420 0.3034 59 0.3144 91 0.3063 $\pm$ 0.0026
Dataset 43 152 0.7011 69 0.6997 86 0.7078 $\pm$ 0.0053
Dataset 44 318 0.4545 73 0.4623 83 0.4563 $\pm$ 0.0041
Dataset 45 48 0.6008 50 0.6047 84 0.6115 $\pm$ 0.0042
Dataset 46 106 0.3603 69 0.3656 88 0.3656 $\pm$ 0.0029
Dataset 47 336 0.5816 58 0.5873 82 0.5893 $\pm$ 0.0034
Dataset 48 133 0.3376 58 0.3392 82 0.3406 $\pm$ 0.0027